# The Jaynes-Cummings model as an autonomous Maxwell demon

### v2

@@ Abstract
Autonomous machines, performing tasks without external control, are an important concept in thermodynamics [1, 2]. Their time evolution derive solely from their initial non-equilibrium condition, akin to automaton. This concept is particularly important in the case of quantum machines [3, 4, 5, 6, 7, 8, 9], where the coherence of microscopic constituents is crucial, as avoiding classical controllers empirically mitigates decoherence. This motivates research into autonomous (or coherent) feedback [10, 11, 12, 13, 14, 15], where a quantum system is able to drive a target system conditionally on the values of a few of its observables, e.g. applied to quantum error correction [16, 17, 18]. When these controllers perform thermodynamic tasks, such as work extraction or cooling, they are called autonomous Maxwell demons [19, 20, 21, 22, 23, 24, 25]. A thorough characterization of these machines is of major fundamental interest, as it is linked to the long-standing question of how the dynamics induced by quantum measurements [26, 27, 28] and the laws of thermodynamics [29, 30, 31] emerge in an entirely quantum framework.
@@
Model – We consider a qubit $Q$ (e.g. a two-level atom) interacting with a harmonic oscillator (hereafter dubbed the cavity $C$) according to the Jaynes-Cummings interaction $\hat{V}_{\text{QC}}=i\hbar g(a\sigma_{+}-a^{\dagger}\sigma_{-})$, where $a$ denotes the bosonic annihilation operator of the cavity, and $\sigma_{-}=|g\rangle\langle e|=\sigma_{+}^{\dagger}$ the lowering operator of the qubit. We focus on resonant conditions, where the qubit and cavity Hamiltonians are $\hat{H}_{Q}=\hbar\omega_{0}\sigma_{+}\sigma_{-}$ and $\hat{H}_{C}=\hbar\omega_{0}a^{\dagger}a$, respectively. We assume that the cavity is initially prepared in a coherently-displaced thermal state $\hat{\rho}_{C}(0)=D(\alpha_{0})\hat{w}_{\beta(0)}D^{\dagger}(\alpha_{0})$, where $\hat{w}_{\beta(0)}=\frac{e^{-\beta(0)\hbar\omega_{0}a^{\dagger}a}}{Z_{C}(0)}$ is a thermal equilibrium state at inverse temperature $\beta(0)$, with $Z_{C}(0)=\text{Tr}\{e^{-\beta(0)\hbar\omega_{0}a^{\dagger}a}\}$. $D(\alpha_{0})=e^{\alpha_{0}a^{\dagger}-\alpha_{0}^{*}a}$ denotes the phase-space displacement operator with amplitude $\alpha_{0}=\sqrt{n_{0}}e^{i\phi_{0}}$. We are interested in the limit $n_{0}\gg\bar{n},1$, with $\bar{n}=(e^{\beta(0)\hbar\omega_{0}}-1)^{-1}$ the initial thermal occupation of the cavity, in which the cavity is expected to induce a classical drive for the qubit.
Figure 1: (a) Entropy variation $\Delta S_{Q}$ of the qubit over a long time scale for different initial qubit states $|\psi_{1}\rangle=|e\rangle\langle e|$, $|\psi_{2}\rangle=\frac{1}{\sqrt{2}}(|e\rangle+e^{i\frac{\pi}{4}}|g\rangle)$, and $|\psi_{3}\rangle=|+_{y}\rangle=\frac{1}{\sqrt{2}}(|e\rangle+i|g\rangle)$. (b) zoom in the early times of (a), in semilog scale. (c) Excited state population $P_{e}=\langle e|\hat{\rho}_{Q}(t)|e\rangle$ (in blue) and coherence $C_{eg}=\langle e|\hat{\rho}_{Q}(t)|g\rangle$ (in orange and red) of the qubit when it is initialized in $\hat{\rho}_{Q}(0)=|e\rangle\langle e|$. The dots correspond to expansions up to second order in $1/n_{0}$ and the solid lines are obtained from the numerical evaluation of the exact expressions, truncating the cavity Hilbert space to $N_{\text{ph}}=500$. Parameters: $n_{0}=100$, $\bar{n}=1$.
@@
We find that the dynamics of the autonomous machine features three successive regimes with very different thermodynamic interpretations, that can be identified from the time-evolution of the qubit’s von Neumann entropy as shown in Fig. 1(a) and (b). We analyze them in the following.
@@
Unitary regime – For times $t\ll g^{-1}$, the cavity and the qubit density operators remain approximately factorized, such that the cavity behaves as a quasi-ideal source of work inducing a classical drive on the qubit. As a consequence, the qubit follows a unitary evolution ruled by the effective Hamiltonian $\hat{H}^{\text{eff}}_{Q}(t)=\text{Tr}_{C}\{\hat{V}_{\text{QC}}\hat{\rho}_{C}(t)\}=i\hbar g(\left\langle a(t)\right\rangle\sigma_{+}-\langle{a^{\dagger}(t)}\rangle\sigma_{-})$, featuring Rabi oscillations with frequency $\Omega\simeq g|\left\langle a(0)\right\rangle|=g\sqrt{n_{0}}$ (see Fig. 1(c) and [32]). To identify thermodynamic evidence of the work source behavior of the cavity, and deviations from it, we plot in Fig. 2 the work $W_{C}$ and the heat $Q_{C}$ provided by the cavity to the qubit, together with the variation of the cavity’s internal energy $\Delta E_{C}$. We see that for large enough values of $n_{0}$, the cavity mostly provides work and almost no entropy is produced during a single Rabi oscillation. We formalize this result by expanding as a power series of $1/n_{0}$ the exact cavity and qubit state in the interaction picture $\hat{\rho}(t)=\hat{U}(t)(\hat{\rho}_{Q}(0)\otimes\hat{\rho}_{C}(0))\hat{U}^{\dagger}(t)$, where the analytical expression of $\hat{U}(t)$ is provided in [32]. For $\hat{\rho}_{Q}(0)=|e\rangle\langle e|$, the qubit population $P_{e}(t)=\langle e|\hat{\rho}_{Q}(t)|e\rangle$ and coherence $C_{eg}(t)=\langle e|\hat{\rho}_{Q}(t)|g\rangle$ read:
@@
where $\theta=2\Omega t=2gt\sqrt{n_{0}}$ stands for the rotation angle in the Bloch sphere. The complete expression of first order corrections $\delta P_{e}(\theta,\bar{n})/n_{0}$ and $\delta C_{eg}(\theta,\bar{n})/n_{0}$ to an ideal Rabi oscillation are given in [32], and decay as $1/n_{0}$ in the classical limit. These first-order truncations show good agreement with exact numerical results over short time-scales of the order of $\Omega^{-1}$, as shown in Fig. 1(c). In addition, our expansion shows that the amount of exchanged heat $Q_{C}$ and the entropy production scale as $1/n_{0}$ (see Fig. 2), confirming that the cavity can be considered as an ideal work source in this regime.
@@
Figure 2: Heat $Q_{C}$ (red), work $W_{C}$ (green) and internal energy variation of the cavity $\Delta E_{C}$ (blue) as a function of $\theta=2gt\sqrt{n_{0}}$, for $\hat{\rho}_{Q}(0)=|e\rangle\langle e|$, $n_{0}=500$ and $\bar{n}=1$. Dots are obtained from the second order expansion in $1/n_{0}$, and exact expressions with truncated cavity Hilbert space are displayed in solid black (not distinguishable). For smaller values of $n_{0}$, the heat provided by the cavity is non-negligible, as shown by the dashed black line for $n_{0}=10$.
@@
Measurement-induced decoherence – It is known from previous studies [33, 34, 35], as well as measurements on Rydberg atoms in cavities [36, 37], that the qubit decoheres, and Rabi oscillations collapse, over a time scale $t_{c}\sim g^{-1}$ (see Fig. 1(c)). This decoherence is due to the evolution of the field towards two orthogonal states, with macroscopically different values of the complex field amplitudes $\left\langle a(t)\right\rangle$ conditioned on the state of the qubit. This evolution can be interpreted as a measurement of the qubit by the cavity  [^1] A more correct terminology here would be a “pre-measurement”. The measurement would be complete only if the measurement outcome is copied in a large number of environmental degrees of freedom [38]. If the cavity and qubit continue to evolve in isolation, revivals of the Rabi oscillations would eventually be observed [35]. However, the measurement analogy remains fully consistent on the time interval considered in this letter.[39, 40].
@@
The measurement basis is set by the eigenbasis of $\hat{H}_{Q}^{\text{eff}}(0)$, and therefore by the initial phase of the field. Explicitly, it reads $\{\cos(\frac{\phi_{0}}{2})|+_{y}\rangle-i\sin(\frac{\phi_{0}}{2})|-_{y}\rangle,-i\sin(\frac{\phi_{0}}{2})|+_{y}\rangle+\cos(\frac{\phi_{0}}{2})|-_{y}\rangle\}$, where |±$|\pm_{y}\rangle$ are the eigenstates of $\sigma_{y}=i(\sigma_{-}-\sigma_{+})$, and corresponds to an axis in the equator of the Bloch sphere. In the case $\phi_{0}=0$, which we consider hereafter, the qubit is measured along the $y$-axis. The total state verifies for $t\lesssim t_{c}={\cal O}(g^{-1})$ [40]:
$\hat{\rho}(t)=\!\!\!\sum_{\nu,\mu=\pm}\rho_{\nu\mu}|\nu_{y}\rangle\langle\mu_{y}|\otimes\hat{\rho}_{C}^{\nu\mu}(t)+{\cal O}\!\left(\frac{1}{n_{0}}\right)\!,\;\;\;$ (2)
@@
where $\rho_{\nu\mu}=\langle\nu_{y}|\rho_{Q}(0)|\mu_{y}\rangle$ are the coefficients of the qubit density operator in the measurement basis and
$\hat{\rho}_{C}^{\nu\mu}(t)\!\!\!$ $=$ $\!\!\!e^{\nu i\chi\delta_{\nu\mu}}D(\sqrt{n_{0}}e^{\nu\frac{igt}{2\sqrt{n_{0}}}})w_{\beta(0)}D^{\dagger}(\sqrt{n_{0}}e^{\mu\frac{igt}{2\sqrt{n_{0}}}}),\;\;\;$ (3)
@@
are conditional operators in the cavity space with $\chi=gt(\sqrt{n_{0}}+1/\sqrt{n_{0}})$ and $\delta_{\nu\mu}$ the Kronecker delta. The diagonal terms $\hat{\rho}_{C}^{\nu\nu}$ are conditional cavity states associated with measurement results $\nu$, and are characterized by field amplitudes
@@
For $t\in[0,t_{c}]$, they fulfill $\alpha_{\nu}(t)=\sqrt{n_{0}}+i\nu\frac{gt}{2}+{\cal O}\left(1/\sqrt{n_{0}}\right)$. The trace of the cross terms $\text{Tr}\{\hat{\rho}_{C}^{+-}(t_{c})\}$ vanishes for $n_{0}\gg\bar{n},1$, which implies maximal qubit-cavity correlations, and hence a complete measurement [41]. In particular, the reduced qubit state verifies $\hat{\rho}_{Q}(t_{c})=\rho_{++}|+_{y}\rangle\langle+_{y}|+\rho_{--}|-_{y}\rangle\langle-_{y}|+{\cal O}(1/n_{0})$, achieving complete dephasing of the Rabi oscillations in the classical limit.

### v4

@@ Abstract
I Introduction
Autonomous machines are a key concept in thermodynamics [1, 2]. Their time evolution solely derives from their initial non-equilibrium condition, akin to an automaton, allowing them to perform tasks without external control. This concept is all the more important in the case of quantum machines [3, 4, 5, 6, 7, 8, 9] where the coherence of microscopic constituents is crucial, as it avoids employing classical controllers which are typically sources of decoherence. The latter observation motivates research into autonomous (or coherent) feedback schemes [10, 11, 12, 13, 14, 15], that is, protocols where a quantum system is able to drive a target system conditionally on the values of a few of its observables, e.g. applied to quantum error correction [16, 17, 18]. When these feedback controllers perform thermodynamic tasks, such as work extraction or cooling, they are called autonomous Maxwell demons [19, 20, 21, 22, 23, 24, 25]. Like in the famous gedanken experiment attributed to Maxwell, the information about the system exploited in the feedback mechanism is a resource, enabling work extraction from a single heat source or cold-to-hot heat transfer in the absence of work, hence apparent violations of the second law if those resources are disregarded. A thorough characterization of these machines is of major fundamental interest, as it is linked to the long-standing question of how the dynamics induced by quantum measurements [26, 27, 28] and the laws of thermodynamics [29, 30, 31] emerge in an entirely quantum framework.
@@ I Introduction
II Thermodynamics of autonomous quantum machines
To interpret the energy exchanges between the cavity and the qubit in terms of work and heat, and deduce the constraints set by the second law of thermodynamics, we rely on the framework introduced by some of us in Ref. [30]. In a nutshell, this framework identifies work with energy exchanges along isoentropic transformation, and heat with energy exchanges proportional to the entropy variations. More precisely, the method applies to an arbitrary set of quantum systems whose evolution is ruled by a total time-independent Hamiltonian
$\hat{H}=\sum_{j}\hat{H}_{j}+\hat{V}$, (1)
where $H_{j}$ stands for the bare Hamiltonian of system $j$, and $\hat{V}$ an arbitrary interaction Hamiltonian. Those systems can include the thermodynamic system of interest (the working body), as well as the energy sources used to power its transformation, treated as quantum systems. Importantly, the role of each system $j$ (e.g. heat source or work source) is not pre-supposed; all systems are a priori hybrid sources of work and heat for the others, and are treated on an equal footing.
The internal energy of system $j$ is defined as $E_{j}(t)=\text{Tr}\{\hat{H}_{j}\hat{\rho}_{j}(t)\}$, computed from its reduced density operator $\hat{\rho}_{j}(t)=\text{Tr}_{j^{\prime}\neq j}\{\hat{\rho}_{\text{tot}}(t)\}$, where $\hat{\rho}_{\text{tot}}(t)$ stands for the density operator of the whole set of systems. Then, the heat provided by system $j$ to the others between times $t=0$ and $t$ is defined as the variation
$Q_{j}(t)=-\Delta E_{j}^{\text{th}},$ (2)
of the thermal energy
$E_{j}^{\text{th}}(t)=\text{Tr}\{\hat{w}_{j}(t)\hat{H}_{j}\}.$ (3)
The latter is the internal energy of the unique thermal state $\hat{w}_{j}(t)=e^{-\beta_{j}(t)\hat{H}_{j}}/Z_{j}(t)$ of system $j$ which has the same von Neumann entropy
$S_{j}(t)=-\text{Tr}\{\hat{\rho}_{j}(t)\log\hat{\rho}_{j}(t)\}$ (4)
as the actual reduced state $\hat{\rho}_{j}(t)$. As thermal states achieve minimal energy at fixed von Neumann entropy, the thermal energy $E_{j}^{\text{th}}(t)\leq E_{j}(t)$ represents the part of the internal energy of system $j$ which carries entropy (or the energy remaining in the system after performing on it the isoentropic transformation extracting the largest amount of energy). For an infinitesimal evolution, the heat verifies
@@ II Thermodynamics of autonomous quantum machines
such that an exchange of heat is necessarily accompanied by a change of entropy of the energy source providing heat. Conversely, the work provided by the system $j$ is defined as the energy exchange associated to the isoentropic part of the evolution of system $j$ and is obtained from:
$W_{j}(t)={\color[rgb]{0,0,0}{-}}\Delta E_{j}-Q_{j}(t).$ (6)
The effective nonequilibrium inverse temperature $\beta_{j}(t)$ is the solution of the equation $S_{j}(t)=-\text{Tr}\{\hat{w}_{j}(t)\log\hat{w}_{j}(t)\}$. It sets constraints on the energy exchanges occurring during the joint evolution of the systems. For instance, for an uncorrelated total initial state of the systems $\hat{\rho}_{\text{tot}}({\color[rgb]{0,0,0}{t_{0}}})=\otimes_{j}\hat{\rho}_{j}({\color[rgb]{0,0,0}{t_{0}}})$, one can express the second law for the transformation of system $j$ under the form of Clausius inequality [30]
$\Delta S_{j}-\sum_{j^{\prime}\neq j}\beta_{j^{\prime}}(0){Q}_{j^{\prime}}(t)\geq 0.$ (7)
From these definitions, a pure source of work, verifying $\Delta E_{j}=W_{j}(t)$, can be identified as a quantum system never building correlations with the others, that is, fulfilling $\hat{\rho}_{\text{tot}}(t)=\hat{\rho}_{j}(t)\otimes\text{Tr}_{j}\{\hat{\rho}_{\text{tot}}(t)\}$. The evolution of system $j$ is then isoentropic, and it effectively behaves as a classical drive on the remaining systems. Such behavior is expected to occur only asymptotically in realistic systems, such as a cavity in the classical limit $n_{0}\gg 1,\bar{n}$, coupled to an atom. The framework presented here quantitatively evaluates deviations from this ideal picture, e.g. residual exchanges of heat associated with the finite perturbation of the cavity state and the correlations generated by the cavity-atom dynamics.
III (Autonomous) Maxwell demons
In a modern formulation [32, 33, 34, 35], Maxwell demons can be defined as entities acquiring information on the quantum system of interest and performing a feedback on it based on this information. If the presence of the demon is not known, and the system only is monitored, the action of the demon results in apparent violations of the second law, such as a decrease of the system’s entropy without an outward heat flow or a cyclic work extraction from a single heat source.
In the historical (non-autonomous) scenario, i.e. Maxwell’s seminal gedanken experiment, the demon is an agent performing a measurement on the system and then storing the result in a memory. It then acts conditionally to this result on the system, allowing for an apparent cyclic work extraction from a single heat source. The paradox is solved by considering the finite work cost to reset the demon’s memory, which exactly compensates the extracted work when the reset process is performed in contact with the same heat source as the work extraction. Alternatively, in non-cyclic transformations, one must consider the information previously acquired by the demon as a resource, which when consumed, plays a role similar to a work expenditure. A similar analysis can be done when the system and the demon are described together as an autonomous system, solely evolving due to the dynamics generated by their joint, time-independent Hamiltonian: in this case, information flows from the system to the demon (which can be seen as the autonomous counterpart of a measurement) and influences the later dynamics induced by the demon on the working system (autonomous feedback). The dynamics of the system under the influence of the demon obeys a modified Clausius inequality, taking into account the information acquired by the demon prior to time $t_{0}$. Focusing on a system S interacting with a demon D, both forming an isolated system, one has [30]:
$\Delta S_{S}-\beta_{D}(0)Q_{D}(t)\geq\Delta I(t).$ (8)
Here $I(t)=S_{S}(t)+S_{D}(t)-S_{SD}(t)\geq 0$ is the quantum mutual information between the demon and the system at time $t$, and $\Delta I(t)=I(t)-I(t_{0})$ its variation between time $t_{0}$ and $t>t_{0}$. $I(t_{0})>0$ quantifies the information acquired by the demon before $t=t_{0}$, and $\Delta I(t)$ the amount of information consumed in the transformation. Comparing with Eq. (7) in the case $j=S$, $j^{\prime}=D$, one can see that the left-hand side can now take negative values provided $I(t)<I_{0}$ (some information is consumed). In other words, information acquired about the system, as quantified by the quantum mutual information, can be seen as a thermodynamic resource that the demon can consume to obtain effects akin to a work expenditure. In the most striking case, the demon provides no work nor heat, and still increases the system’s free energy, or decreases its entropy [35].
IV Physical Model
From now on, we consider the case of a bi-partite system $j,j^{\prime}=\text{Q},\text{C}$, where Q stands for a qubit (e.g. a two-level atom) and C a harmonic oscillator (hereafter dubbed the cavity), coupled via the Jaynes-Cummings interaction
$\hat{V}_{\text{QC}}=i\hbar g(a\sigma_{+}-a^{\dagger}\sigma_{-}),$ (9)
where $a$ denotes the bosonic annihilation operator of the cavity, and $\sigma_{-}=|g\rangle\langle e|=\sigma_{+}^{\dagger}$ the lowering operator of the qubit. We focus on resonant conditions, where the qubit and cavity Hamiltonians are $\hat{H}_{\text{Q}}=\hbar\omega_{0}\sigma_{+}\sigma_{-}$ and $\hat{H}_{\text{C}}=\hbar\omega_{0}a^{\dagger}a$, respectively. We assume that the cavity is initially prepared in a coherently-displaced thermal state
$\hat{\rho}_{\text{C}}(0)=D(\alpha_{0})\hat{w}_{\beta_{C}(0)}D^{\dagger}(\alpha_{0}),$ (10)
where $\hat{w}_{\beta_{C}(0)}=\frac{e^{-\beta_{C}(0)\hbar\omega_{0}a^{\dagger}a}}{Z_{\text{C}}(0)}$ is a thermal equilibrium state at inverse temperature $\beta_{C}(0)$, with $Z_{\text{C}}(0)=\text{Tr}\{e^{-\beta_{C}(0)\hbar\omega_{0}a^{\dagger}a}\}$ and
$\bar{n}=(e^{\beta_{C}(0)\hbar\omega_{0}}-1)^{-1}$ (11)
is the initial number of thermal excitations in the oscillator. $D(\alpha_{0})=e^{\alpha_{0}a^{\dagger}-\alpha_{0}^{*}a}$ denotes the phase-space displacement operator with amplitude
$\alpha_{0}=\sqrt{n_{0}}e^{i\phi_{0}}.$ (12)
We are interested in the limit $n_{0}\gg\bar{n},1$, in which the cavity is expected to induce a classical drive for the qubit.
It is useful for the following to emphasize that the model verifies $\hat{H}_{\text{C}}+\hat{H}_{\text{Q}},\hat{V}_{\text{QC}}]=0$, such that the sum of the qubit and cavity internal energies $E_{C}(t)+E_{Q}(t)$ is a constant of motion, and that over any time interval
$\Delta E_{\text{C}}=-\Delta E_{\text{Q}}.$ (13)
Figure 1: (a) Entropy variation $\Delta S_{\text{Q}}$ of the qubit over a long time scale for different initial qubit states $|\psi_{1}\rangle=|e\rangle\langle e|$, $|\psi_{2}\rangle=\frac{1}{\sqrt{2}}(|e\rangle+e^{i\frac{\pi}{4}}|g\rangle)$, and $|\psi_{3}\rangle=|+_{y}\rangle=\frac{1}{\sqrt{2}}(|e\rangle+i|g\rangle)$. (b) zoom in the early times of (a), in semilog scale. (c) Excited state population $P_{e}=\langle e|\hat{\rho}_{\text{Q}}(t)|e\rangle$ (in blue) and coherence $C_{eg}=\langle e|\hat{\rho}_{\text{Q}}(t)|g\rangle$ (in orange and red) of the qubit when it is initialized in $\hat{\rho}_{\text{Q}}(0)=|e\rangle\langle e|$. The dots correspond to expansions up to second order in $1/n_{0}$ and the solid lines are obtained from the numerical evaluation of the exact expressions, truncating the cavity Hilbert space to $N_{\text{ph}}=500$. Parameters: $n_{0}=100$, $\bar{n}=1$.
Applying the framework introduced in Secs. II-III to this model, we find that the dynamics of the autonomous machine features three successive temporal regimes, in which, respectively, the cavity behaves as a quasi-ideal work source (unitary regime), then acquires information about the system (measurement regime), and finally performs autonomous feedback on the qubit. Those three regimes can be identified from the time-evolution of the qubit’s von Neumann entropy as shown in Fig. 1(a) and (b). We analyze them in the following sections, and show the cavity behaves as an autonomous Maxwell demon.
V Unitary regime
For times $t\ll g^{-1}$, the cavity and the qubit density operators remain approximately factorized, such that the cavity behaves as a quasi-ideal source of work inducing a classical drive on the qubit. As a consequence, the qubit follows a unitary evolution ruled by the effective Hamiltonian
$\hat{H}^{\text{eff}}_{\text{Q}}(t)$ $=\text{Tr}_{\text{C}}\{\hat{V}_{\text{QC}}\hat{\rho}_{\text{C}}(t)\}$ (14)
$=i\hbar g(\left\langle a(t)\right\rangle\sigma_{+}-\langle{a^{\dagger}(t)}\rangle\sigma_{-}).$
@@ V Unitary regime
The latter induces Rabi oscillations with an approximately constant frequency $\Omega\simeq g|\!\left\langle a(0)\right\rangle\!|=g\sqrt{n_{0}}$ (see Fig. 1(c) and Appendix B). To identify thermodynamic evidence of the work source behavior of the cavity, and deviations from it, we plot in Fig. 2 the work $W_{\text{C}}$ and the heat $Q_{\text{C}}$ provided by the cavity to the qubit, together with the variation of the cavity’s internal energy $\Delta E_{\text{C}}$. We see that for large enough values of $n_{0}$, the cavity mostly provides work and almost no entropy is produced during a single Rabi oscillation. We formalize this result by expanding as a power series of $1/n_{0}$ the exact cavity and qubit state in the interaction picture
$\hat{\rho}(t)=\hat{U}(t)(\hat{\rho}_{\text{Q}}(0)\otimes\hat{\rho}_{\text{C}}(0))\hat{U}^{\dagger}(t),$ (15)
where the analytical expression of $\hat{U}(t)$ is provided in Appendix A. For $\hat{\rho}_{\text{Q}}(0)=|e\rangle\langle e|$, the qubit population $P_{e}(t)=\langle e|\hat{\rho}_{\text{Q}}(t)|e\rangle$ and coherence $C_{eg}(t)=\langle e|\hat{\rho}_{\text{Q}}(t)|g\rangle$ read:
@@
where $\theta=2\Omega t=2gt\sqrt{n_{0}}$ corresponds to the rotation angle in the Bloch sphere along an axis of the equator with azimuthal angle $\phi_{0}$. The first order corrections to an ideal unitary dynamics are given by
$\delta P_{e}\!\left(\theta,\bar{n}\right)$ $=$ $\!\!\!\!\!\!-\frac{\theta}{16}\left(\left(1+2\bar{n}\right)\theta\cos{\theta}+\left(3+2\bar{n}\right)\sin{\theta}\right)$ (17)
$\delta C_{eg}\!\left(\theta,\bar{n}\right)$ $=$ $\!\!\!\!\!\!\frac{-2\theta\!\!+\!\!\left(3\!+\!2\bar{n}\right)\!\theta\!\cos\!\theta\!-\!\left(1\!+\!2\bar{n}\right)\!\!\left(\!1\!+\!\theta^{2}\right)\!\sin\!\theta}{16}.$
These first-order truncations show good agreement with exact numerical results over time-scales much shorter than $g^{-1}$, as shown in Fig. 1(c). In addition, our expansion shows that the amount of exchanged heat
$Q_{\text{C}}\!\left(t\right)\!=\!-\frac{\hbar\omega_{0}}{16n_{0}}$ $\!\left(\theta^{2}\!+\!\sin^{2}{\theta}\!+\!2\left(1+2\bar{n}\right)\!\theta\sin\!{\theta}\right)$ (18)
$+\!{\cal O}\left(\frac{1}{n_{0}^{2}}\right)$
scales as $1/n_{0}$ (see Fig. 2), confirming that the cavity tends to an ideal work source behavior in the limit $n_{0}\to\infty$. Similarly, entropy production associated with the qubit’s transformation is found to vanish for large $n_{0}$ (see Appendix B).
@@
Figure 2: Heat $Q_{\text{C}}$ (red), work $W_{\text{C}}$ (green) and internal energy variation of the cavity $\Delta E_{\text{C}}$ (blue) as a function of $\theta=2gt\sqrt{n_{0}}$, for $\hat{\rho}_{\text{Q}}(0)=|e\rangle\langle e|$, $n_{0}=500$ and $\bar{n}=1$. Dots are obtained from the second order expansion in $1/n_{0}$, and exact expressions with truncated cavity Hilbert space are displayed in solid black (not distinguishable). For smaller values of $n_{0}$, the heat provided by the cavity is non-negligible, as shown by the dashed black line for $n_{0}=10$.
VI Measurement-induced decoherence
@@
It is known from previous studies [36, 37, 38], as well as measurements on Rydberg atoms in cavities [39, 40], that the qubit decoheres, and Rabi oscillations collapse, over a time scale $t_{c}\sim g^{-1}$ (see Fig. 1(c)). This decoherence is due to the evolution of the field towards two orthogonal states, with macroscopically different values of the complex field amplitudes $\left\langle a(t)\right\rangle$ conditioned on the state of the qubit. This evolution can be interpreted as an autonomous “measurement” of the qubit by the cavity  [^1] A more correct terminology here would be a “pre-measurement”. The measurement process would be complete, and the result would truly become an objective fact, only if it were copied multiple times into the state of a large number of different environmental degrees of freedom [41]. If the cavity and qubit continue to evolve in isolation, revivals of the Rabi oscillations would eventually be observed [38]. However, the measurement analogy remains fully consistent on the time interval considered in this article. [42, 43].
@@
The measurement basis is set by the eigenbasis of $\hat{H}_{\text{Q}}^{\text{eff}}(0)$, and therefore by the initial phase of the field. Explicitly, it reads $\{\cos(\frac{\phi_{0}}{2})|+_{y}\rangle-i\sin(\frac{\phi_{0}}{2})|-_{y}\rangle,-i\sin(\frac{\phi_{0}}{2})|+_{y}\rangle+\cos(\frac{\phi_{0}}{2})|-_{y}\rangle\}$, where $|\pm_{y}\rangle$ are the eigenstates of $\sigma_{y}=i(\sigma_{-}-\sigma_{+})$. It therefore corresponds to a measurement along the rotation axis of the unitary regime, which lies in the equator of the Bloch sphere. In the case $\phi_{0}=0$, which we consider hereafter, the qubit is measured along the $y$-axis. The total state verifies for $t\lesssim t_{c}={\cal O}(g^{-1})$ (see Appendix C.1):
$\hat{\rho}(t)=\!\!\!\sum_{\nu,\mu=\pm}\rho_{\nu\mu}|\nu_{y}\rangle\langle\mu_{y}|\otimes\hat{\rho}_{\text{C}}^{\nu\mu}(t)+{\cal O}\!\left(\frac{1}{n_{0}}\right)\!,\;\;\;$ (19)
@@
where $\rho_{\nu\mu}=\langle\nu_{y}|\rho_{\text{Q}}(0)|\mu_{y}\rangle$ are the coefficients of the qubit density operator in the measurement basis and
$\hat{\rho}_{\text{C}}^{\nu\mu}(t)\!\!\!$ $=$ $\!\!\!e^{\nu i\chi\delta_{\nu\mu}}D(\sqrt{n_{0}}e^{\nu\frac{igt}{2\sqrt{n_{0}}}})\hat{w}_{\beta_{C}(0)}D^{\dagger}(\sqrt{n_{0}}e^{\mu\frac{igt}{2\sqrt{n_{0}}}}),\;\;\;\;$ (20)
@@
are conditional operators in the cavity space with $\chi=gt(\sqrt{n_{0}}+1/\sqrt{n_{0}})$ and $\delta_{\nu\mu}$ the Kronecker delta. The diagonal terms $\hat{\rho}_{\text{C}}^{\nu\nu}$ are conditional cavity states associated with measurement results $\nu=\pm$, and are characterized by field amplitudes
@@
For $t\in[0,t_{c}]$, they fulfill
$\alpha_{\nu}(t)=\sqrt{n_{0}}+i\nu\frac{gt}{2}+{\cal O}\left(1/\sqrt{n_{0}}\right),$ (22)
and therefore differ by the sign of their purely imaginary component. For $t\sim g^{-1}$, those two field amplitude become fully distinguishable (correspond to orthogonal cavity states). The trace of the cross terms $\text{Tr}\{\hat{\rho}_{\text{C}}^{+-}(t_{c})\}$ vanishes for $n_{0}\gg\bar{n},1$, which implies maximal qubit-cavity correlations, and hence a complete measurement [44]. In particular, the reduced qubit state then verifies
$\hat{\rho}_{\text{Q}}(t_{c})=\rho_{++}|+_{y}\rangle\langle+_{y}|+\rho_{--}|-_{y}\rangle\langle-_{y}|+{\cal O}(\frac{1}{n_{0}}),$ (23)
achieving complete dephasing of the Rabi oscillations in the classical limit.
VII Autonomous feedback and qubit purification